OpenStudy (anonymous):

Find the derivative of f(x) =-5/x at x=11

OpenStudy (midhun.madhu1987):

Derivative of 1/x is -1/x^2

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (midhun.madhu1987):

can u try?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

Right you should definitely try this out first \[\large f(x) = \frac{-5}{x}\] We know 5 is just multiplying whatever answer we get, so we can factor that out for now \[\large 5(f(x)) = \frac{1}{x}\] So as stated above, the derivative of 1/x = -1/x^2 so \[\large 5 \times -\frac{1}{x^2} =?\]

OpenStudy (johnweldon1993):

And so all you do after this, is plug in 11 for 'x' and see what result you get

OpenStudy (johnweldon1993):

Sorry put an extra '-' sign up there Since we did the derivative of -1/x we should have canceled out the '-' sign so \[\large \frac{5}{x^2}\] Now you just plug in 11 for 'x'

OpenStudy (anonymous):

so my answer is 5/121?